# An image-computable model of speeded decision-making

### v1

@@ Discussion
In summary, the VAM is a probabilistic model of psychophysical data that captures how raw sensory inputs are transformed into the abstract representations that guide decisions. Raw (pixel-space) visual stimuli are processed by a biologically-plausible neural network model of vision that outputs the parameters of a traditional decision-making model. Each VAM is fitted to data from a single participant, a feature that allowed us to study how individual differences in behavior emerge from differences in the "brains" of the models. To this end, we found that models with smaller congruency effects had more orthogonal representations for task-relevant and irrelevant information. While we chose to use a CNN to model visual processing, we note that the VAM is not limited to this choice: other sensory encoding models, such as those based on transformer architectures [15], can be readily swapped in to replace the CNN with minimal changes to the underlying VAM implementation. Similarly, the LBA decision-making model we employed is easily replaced with other decision-making models that have a closed-form or easily approximated likelihood, such as the diffusion-decision model and leaky competing accumulator model [60, 48, 80, 43]. In this way, the VAM provides a general probabilistic framework for jointly fitting interpretable decision-making models and expressive neural network architectures of sensory processing with psychophysical data.
@@ Methods
We used deidentified Lost in Migration gameplay data from 75 Lumosity users (participants) to train the models included in this paper. The mean (SD) age of this sample at the time of signup was 56.4 (18.6) years; 46.7% identified as female, 48.0% identified as male, and 5.3% did not report their gender. All analysis and modeling was done retrospectively on preexisting Lumosity data that were collected during the normal use of the Lumosity program (at home). All participants consented to the use and disclosure of their deidentified Lumosity data for any purpose as laid out in the Lumosity Privacy Policy (www.lumosity.com/legal/privacy_policy). No statistical methods were used to predetermine sample sizes, though our sample sizes are comparable to or larger than those used in related modeling work [8, 82].
@@
Linear ballistic accumulator (LBA) model. The decision-making component of our modeling framework is the LBA model [8], tailored to Lost in Migration. Evidence for each of the four possible response directions is accumulated linearly and independently at a response-specific drift rate $d_{k}$. Commitment to a decision occurs when one of the accumulators reaches a fixed threshold parameter $b$. The RT on a given trial is the duration of this evidence accumulation process plus a constant non-decision time parameter $t_{0}$ that includes both sensory processing and motor execution. Response variability comes from two sources. First, on each trial, the drift rate $d_{k}$ for each response $k\in\{1,2,3,4\}$ is sampled independently from a Gaussian distribution with mean $v_{k}$ and a common SD $s$. We fix $s$ to 1 for all models to ensure that the LBA parameters are identifiable [11, 25]. Second, the initial evidence for each accumulator is sampled independently from a uniform distribution on the interval [0, $A$], where $A$ is a model parameter.
@@
where the drift rate means $v^{(i)}$ depend on a CNN with parameters $\boldsymbol{\zeta}$ (described below). The factor on the left of Equation (1) is the likelihood of the LBA model, which has a closed-form solution [8]. The factor on the right is the prior distribution over the LBA parameters, specified as a standard multivariate Gaussian $p(\boldsymbol{\theta})\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. We express the joint density more compactly as:
@@
To fit the VAM, i.e. learn the posterior distribution over the LBA parameters $p(\boldsymbol{\theta}|\mathbf{x})$ and simultaneously optimize the CNN parameters $\boldsymbol{\zeta}$, we apply automatic differentiation variational inference (ADVI) [38]. Rather than attempting to sample from the true posterior directly as in Markov chain Monte Carlo (MCMC) methods, variational inference introduces an approximate posterior density $q(\boldsymbol{\theta};\boldsymbol{\phi})$ and minimizes the Kullback-Leibler (KL) divergence from $p(\boldsymbol{\theta}|\mathbf{x})$ to $q(\boldsymbol{\theta};\boldsymbol{\phi})$ by optimizing $\boldsymbol{\phi}$. Here, we specify the approximate posterior $q(\boldsymbol{\theta};\boldsymbol{\phi})$ as a multivariate Gaussian with mean $\boldsymbol{\mu}$ and unconstrained covariance matrix $\boldsymbol{\Sigma}$ (thus $\boldsymbol{\phi}=\{\boldsymbol{\mu},\boldsymbol{\Sigma}\}$).
We use variational inference since it scales well to large datasets and can handle complicated models (such as the VAM), in contrast to MCMC methods. Leveraging automatic differentiation software (JAX [6]), we automate the calculation of the derivatives of the variational objective (described below) with respect to both the CNN parameters $\boldsymbol{\zeta}$ and variational parameters $\boldsymbol{\phi}$.
Variable transformations. Note that the LBA parameters $\boldsymbol{\theta}=\{b$, $A$, $t_{0}$} are restricted to be non-negative, while $\boldsymbol{\mu}\in\mathbb{R}^{3}$ and $\boldsymbol{\Sigma}$ is restricted to be positive semidefinite. However, to optimize the variational objective (defined below), we require that $\boldsymbol{\phi}$ and $\boldsymbol{\theta}$ have the same support [38]. To achieve this, we transform $\boldsymbol{\phi}$ and $\boldsymbol{\theta}$ to both have support on the real line. Specifically, we reparameterize $\boldsymbol{\Sigma}$ using the Cholesky decomposition as $\boldsymbol{\Sigma}=\mathbf{LL}^{\intercal}$, where $\mathbf{L}$ is a lower-triangular matrix with entries in $\mathbb{R}$ (so that now $\boldsymbol{\phi}=\{\boldsymbol{\mu},\mathbf{L}\}$). For the LBA parameters, in addition to mapping them to the real line, we must also enforce the constraint that the threshold $b$ is always greater than the parameter $A$ controlling the range of the initial evidence distribution. We enforce this by defining $\tilde{b}=b-A$ and taking the log of $\tilde{b}$, $A$, and $t_{0}$ to map them to the real line [11]. We define the transformed parameters as $\boldsymbol{\theta}^{\ast}=\{b^{\ast},A^{\ast},t_{0}^{\ast}\}=\{\log{(b-A)},\log{A},\log{t_{0}}\}$ and the transformation that maps $\boldsymbol{\theta}$ to $\boldsymbol{\theta}^{\ast}$ as $\mathit{T}$.
@@
We will apply stochastic gradient ascent to maximize the ELBO objective function given by Equation (4), using Monte Carlo (MC) estimates of the expectation and AD to calculate gradients with respect to $\boldsymbol{\phi}$ and $\boldsymbol{\zeta}$. However, the gradients of the ELBO with respect to $\boldsymbol{\phi}$ cannot be calculated directly by AD, since the expectation in Equation (4) is taken with respect to $q(\boldsymbol{\theta}^{\ast};\boldsymbol{\phi})$, which depends on $\boldsymbol{\phi}$. We work around this by applying the reparameterization trick [61, 33], which expresses $\boldsymbol{\theta}^{\ast}$ as a differentiable transformation of an auxiliary noise variable $\boldsymbol{\epsilon}\sim p(\boldsymbol{\epsilon})$, where $p(\boldsymbol{\epsilon})$ does not depend on $\boldsymbol{\phi}$. In particular, we set $p(\boldsymbol{\epsilon})$ to be a standard multivariate Gaussian, $p(\boldsymbol{\epsilon})=\mathcal{N}(\boldsymbol{\epsilon};\mathbf{0},\textbf{I})$, and define the transformation $\widetilde{\boldsymbol{\theta}}^{\ast}=\mathbf{L}\boldsymbol{\epsilon}+\boldsymbol{\mu}$, where $\boldsymbol{\mu}$ and $\mathbf{L}$ are the mean and covariance parameters of the Gaussian approximating density $q(\boldsymbol{\theta}^{\ast};\boldsymbol{\phi})$.
Now we estimate the expectation in Equation (4) with MC samples from $p(\boldsymbol{\epsilon})$. Since the expectation no longer depends on $\boldsymbol{\phi}$, we can use AD directly on these MC estimates to obtain unbiased gradients of the ELBO. We estimate the ELBO for each data point using independent MC samples, since this reduces the variance of the gradients relative to using the same set of MC samples for a given batch of data [35]. Thus the ELBO objective for the $i$th data point is estimated by:
@@
VAM inference algorithm and other training details. The VAM training/inference algorithm is summarized in Algorithm 1. The size of the training set was 16,250 samples (65% of the total dataset) for each model. An additional validation set of 3,750 samples (15%) was used to monitor model training. The remaining 5,000 samples (20%) were used to evaluate model performance (the holdout set). We set the batch size $M$ to 256 and the number of MC samples used to estimate the ELBO $L$ to 10. We used the Adam optimizer [34] with the following hyperparameters for model training: learning rate $=1e-3$, $\beta_{1}=0.9$, and $\beta_{2}=0.999$. The same random seed was used to initialize the parameters of all models. Models were trained using a single NVIDIA GeForce RTX 3060 Ti GPU. Each model took approximately one hour to train.
@@
Convolutional neural network (CNN). We used the same seven-layer CNN architecture for all models (six convolutional layers followed by one fully-connected hidden layer). The number of channels/units used in the seven layers was as follows: 64, 64, 128, 128, 128, 256, 1024. The output layer (after the fully-connected layer) has four channels, one for each drift rate. Each convolutional layer used a 3$\times$3 pixel kernel (stride 1, same padding). Each convolutional layer was followed by a ReLU nonlinearity, instance normalization [79], and a max-pooling layer (2$\times$2 pixel window, stride 2), in that order. The fully-connected hidden layer was followed by a ReLU nonlinearity and a dropout layer (dropout rate set to 0.5).
To speed up model training, the first two convolutional layers were initialized with the parameters from a larger CNN trained on an image classification task. Specifically, the pretrained model used a 16-layer VGG architecture and was trained on the ImageNet dataset [12, 70]. These first two layers were trainable (i.e. not fixed to their initial values). The weights of the other convolutional layers, fully-connected layer and output layer were initialized randomly using the LeCun normal initializer [36]; the biases were initialized with zeros.
@@
We used data augmentation techniques commonly used in other image classification training paradigms to improve the generalization ability of the models. Specifically, for each batch of training data, each image was independently and randomly translated by a small amount. The size of the translation (in pixels) was drawn from a uniform distribution on the interval [0,1] (vertical) and [0,2] (horizontal), rounded to the nearest pixel (horizontal and vertical translations were sampled independently). We also applied a variation of a random elastic image deformation [68] as implemented in the Augmax Python package [26]. Specifically, we used the Warp transformation in Augmax and set the strength parameter to 3 and the coarseness parameter to 32. The transformation was applied to each image independently with a probability of 0.75.
@@
The task-optimized CNN models were trained by minimizing the cross-entropy loss function, where the correct label was defined as the direction of the target bird. Other than the loss function, all aspects of the training process were the same as those used for the VAM (CNN architecture, optimizer settings, initialization, etc.). For each VAM we trained, we trained one task-optimized model using the same training data ($n$ = 75 task-optimized models).
Models trained on separate RT quantiles
For the conditional accuracy function and delta plot results presented in Fig. S4C-D, we trained models on data from $n$ = 10 randomly selected participants using a modified training paradigm. For each participant, we calculated the quintiles (i.e. the 0.2, 0.4, 0.6, and 0.8 quantiles) of their RT distribution and trained a separate VAM on the trials within each RT quantile bin (five models per participant). As for the VAMs used in the original training paradigm, 65% of the data within each RT quantile bin was used for training, 15% was used to monitor model training, and the remaining 20% was designated as holdout data used to evaluate model performance. We generated RTs and choices from the holdout stimuli for each of the five models from a given participant, then combined these outputs to compare with the participant’s data.
For a subset of attempted training runs (7/57 = 12.3%, 7 different participants), the model failed to converge (near chance accuracy for 6/7 models, 69% accuracy for 1/7 models). In these cases, we trained a new model from the same participant and RT bin, but used a different random seed to initialize the model parameters and partition the training set vs. holdout set (all such retrained models converged successfully). We did not observe a clear association between training success rate and RT quantile bin (number of failed models in RT quantile bins 1-5: 0, 2, 1, 2, 2).
@@
We analyzed all of the VAMs after 12,800 parameter update steps (corresponding to the 200th training epoch), by which point training had converged. We analyzed all of the task-optimized models after 1,600 parameter update steps (the 25th training epoch) since these models converged much more quickly. The VAMs trained on data from a single RT quantile bin were all analyzed after 6,500 parameter update steps (the 500th training epoch, given the smaller training set size) since these models also converged more quickly than those trained with the entire dataset. All analyses were conducted using a holdout set of $n$ = 5000 LIM stimuli/RTs/choices (i.e., data that was not used to train the models). To generate RTs and choices from the trained models, the LIM stimuli from the holdout set were provided as inputs to the CNN, which output mean drift rates for each stimulus. We denote the drift rate mean for the $k^{th}$ accumulator on trial $i$ as $v_{k}^{(i)}$. We used the LBA model to sample RTs and choices using these mean drift rates, where the drift rate for the $k^{th}$ accumulator on trial $i$, $d_{k}^{(i)}$, is sampled from a normal distribution with mean $v_{k}^{(i)}$ and SD = 1. For a very small fraction of trials (mean $\pm$ s.e.m. percent excluded trials: 0.080 $\pm$ 0.011%, $n$ = 75 models), all of the sampled drift rates were negative, and thus the RT and choice were undefined. These trials were excluded from all analyses.
@@
The models/participants included in the analysis of how RT varied with stimulus layout and horizontal/vertical position were those that exhibited significant modulation of RT by one or more of these stimulus features. Significant modulation was determined by running an ANOVA on the RTs in each stimulus feature bin (e.g. for each layout or each horizontal position bin), using a threshold $p$-value of 0.05. For horizontal stimulus position, we used bins of width = 50 pixels in the original 640$\times$480 pixel window space, except for the leftmost and rightmost bins which had width = 60 pixels. For vertical stimulus position, we used bins of width = 25 pixels.
@@
CNN features. The model features used in the analyses of CNN representations were derived from the unit activations in each layer elicited by the holdout image set. The activations were processed just after the ReLU nonlinearity. The features in the convolutional layers are defined as the maximum value of each channel (i.e., the maximum was calculated across the spatial dimensions) [27], yielding a $N\times K_{l}$ feature matrix, where $N$ is the number of images in the holdout set (5000) and $K_{l}$ is the number of active channels in layer $l$. For the fully-connected layer, the features were defined as the $N\times K_{l}$ matrix of activations.
@@
Only incongruent trials were used for all of the analyses involving the CNN features described below, for the following reasons. For the selectivity and tolerance analyses in which we decoded target or flanker direction from the model features in each layer, described in more detail below, note that the target and flanker directions are by definition the same on congruent trials. As such, a classifier trained to decode target direction from congruent trials could achieve perfect accuracy using information in the model features attributable to the flankers, rather than targets. Using only incongruent trials ensured that such cross-contamination could not occur. We used only incongruent trials for all of the other analyses of the model features simply for convenience.
@@
Stimulus feature decoding. To assess how well particular stimulus features could be decoded from the activity in each layer, we trained a linear SVM to classify the values of that stimulus feature using the $N\times K_{l}$ activation (feature) matrix in each layer. The activation matrix was standardized before training the classifiers, such that each column had zero mean and unit variance. Horizontal and vertical stimulus positions were discretized to enable classification using the same bins as were used in the behavioral analyses.
@@
The classification task was done in a standard "one-vs-rest" setting: for each value of a given stimulus feature, one sub-classifier was trained on a binary classification task with the chosen value as one class and all other values as the other class, yielding one classifier for each value of the stimulus feature (e.g. four for target direction, seven for stimulus layout). To determine the decision of the combined classifier for a given image, we generated predictions from each sub-classifier and assigned the decision to the sub-classifier with the largest (most confident) prediction. We assessed the overall decoding accuracy of each SVM on a separate test image set. Note that both the training set and test set for the SVMs were subsets of the holdout image set (i.e., the images that were not used to train the VAMs). The SVMs were trained using the LinearSVC model in the scikit-learn Python package [54] with the squared hinge loss function and L2 regularization with the penalty parameter C set to 1.0 (the default settings).
@@
Tolerance. To assess the tolerance of model representations to variation in a given stimulus feature (flanker direction, stimulus layout, horizontal position, and vertical position), we trained a linear SVM to classify target direction using stimuli with the chosen stimulus feature fixed to one value (the training context), and assessed the generalization performance of the SVM on stimuli that contained all other values of that stimulus feature (the generalization context). For example, to assess tolerance to stimulus layout, we trained one SVM to classify target direction using stimuli with the vertical line layout, and assessed the generalization performance of that classifier on stimuli with the six other layouts. We trained one such SVM for each of the seven stimulus layouts, and averaged the generalization performance across these seven SVMs to derive the overall generalization performance measure for a given model and network layer. Other details of the classifiers were the same as those used for the decoding analyses described above.

### v2

@@ Conclusion
The VAM is a probabilistic model of psychophysical data that captures how raw sensory inputs are transformed into the abstract representations that guide decisions. Raw (pixel-space) visual stimuli are processed by a biologically-plausible neural network model of vision that outputs the parameters of a traditional decision-making model. Each VAM is fitted to data from a single participant, a feature that allowed us to study how individual differences in behavior emerge from differences in the "brains" of the models. To this end, we found that models with smaller congruency effects had more orthogonal representations for task-relevant and irrelevant information. While we chose to use a CNN to model visual processing, we note that the VAM is not limited to this choice: other sensory encoding models, such as those based on transformer architectures [dosovitskiy2021ViT], can be readily swapped in to replace the CNN with minimal changes to the underlying VAM implementation. Similarly, the LBA decision-making model we employed is easily replaced with other decision-making models that have a closed-form or easily approximated likelihood, such as the diffusion-decision model and leaky competing accumulator model [ratcliff19982choice, navarro2009_first_passage, usher2001, lo2021_LCA_likelihood]. In this way, the VAM provides a general probabilistic framework for jointly fitting interpretable decision-making models and expressive neural network architectures of sensory processing with psychophysical data.
@@ Methods
We used deidentified Lost in Migration gameplay data from 75 Lumosity users (participants) to train the models included in this paper. The mean (SD) age of this sample at the time of signup was 56.4 (18.6) years; 46.7% identified as female, 48.0% identified as male, and 5.3% did not report their gender. All analysis and modeling was done retrospectively on preexisting Lumosity data that were collected during the normal use of the Lumosity program (at home). All participants consented to the use and disclosure of their deidentified Lumosity data for any purpose as laid out in the Lumosity Privacy Policy (www.lumosity.com/legal/privacy_policy). No statistical methods were used to predetermine sample sizes, though our sample sizes are comparable to or larger than those used in related modeling work [brown2008, white2011].
@@
Linear ballistic accumulator (LBA) model. The decision-making component of our modeling framework is the LBA model [brown2008], tailored to Lost in Migration. Evidence for each of the four possible response directions is accumulated linearly and independently at a response-specific drift rate $d_{k}$. Commitment to a decision occurs when one of the accumulators reaches a fixed threshold parameter $b$. The RT on a given trial is the duration of this evidence accumulation process plus a constant non-decision time parameter $t_{0}$ that includes both sensory processing and motor execution. Response variability comes from two sources. First, on each trial, the drift rate $d_{k}$ for each response $k\in\{1,2,3,4\}$ is sampled independently from a Gaussian distribution with mean $v_{k}$ and a common SD $s$. We fix $s$ to 1 for all models to ensure that the LBA parameters are identifiable [dao2022, gunawan2020]. Second, the initial evidence for each accumulator is sampled independently from a uniform distribution on the interval [0, $A$], where $A$ is a model parameter.
@@
where the drift rate means $v^{(i)}$ depend on a CNN with parameters $\boldsymbol{\zeta}$ (described below). The factor on the left of Equation (1) is the likelihood of the LBA model, which has a closed-form solution [brown2008]. The factor on the right is the prior distribution over the LBA parameters, specified as a standard multivariate Gaussian $p(\boldsymbol{\theta})\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. We express the joint density more compactly as:
@@
To fit the VAM, i.e., learn the posterior distribution over the LBA parameters $p(\boldsymbol{\theta}|\mathbf{x})$ and simultaneously optimize the CNN parameters $\boldsymbol{\zeta}$, we apply automatic differentiation variational inference (ADVI) [kucukelbir2017]. Rather than attempting to sample from the true posterior directly as in Markov chain Monte Carlo (MCMC) methods, variational inference introduces an approximate posterior density $q(\boldsymbol{\theta};\boldsymbol{\phi})$ and minimizes the Kullback-Leibler (KL) divergence from $p(\boldsymbol{\theta}|\mathbf{x})$ to $q(\boldsymbol{\theta};\boldsymbol{\phi})$ by optimizing $\boldsymbol{\phi}$. Here, we specify the approximate posterior $q(\boldsymbol{\theta};\boldsymbol{\phi})$ as a multivariate Gaussian with mean $\boldsymbol{\mu}$ and unconstrained covariance matrix $\boldsymbol{\Sigma}$ (thus $\boldsymbol{\phi}=\{\boldsymbol{\mu},\boldsymbol{\Sigma}\}$).
We use variational inference since it scales well to large datasets and can handle complicated models (such as the VAM), in contrast to MCMC methods. Leveraging automatic differentiation software (JAX [jax2018github]), we automate the calculation of the derivatives of the variational objective (described below) with respect to both the CNN parameters $\boldsymbol{\zeta}$ and variational parameters $\boldsymbol{\phi}$.
Variable transformations. Note that the LBA parameters $\boldsymbol{\theta}=\{b$, $A$, $t_{0}$} are restricted to be non-negative, while $\boldsymbol{\mu}\in\mathbb{R}^{3}$ and $\boldsymbol{\Sigma}$ is restricted to be positive semidefinite. However, to optimize the variational objective (defined below), we require that $\boldsymbol{\phi}$ and $\boldsymbol{\theta}$ have the same support [kucukelbir2017]. To achieve this, we transform $\boldsymbol{\phi}$ and $\boldsymbol{\theta}$ to both have support on the real line. Specifically, we reparameterize $\boldsymbol{\Sigma}$ using the Cholesky decomposition as $\boldsymbol{\Sigma}=\mathbf{LL}^{\intercal}$, where $\mathbf{L}$ is a lower-triangular matrix with entries in $\mathbb{R}$ (so that now $\boldsymbol{\phi}=\{\boldsymbol{\mu},\mathbf{L}\}$). For the LBA parameters, in addition to mapping them to the real line, we must also enforce the constraint that the threshold $b$ is always greater than the parameter $A$ controlling the range of the initial evidence distribution. We enforce this by defining $\tilde{b}=b-A$ and taking the log of $\tilde{b}$, $A$, and $t_{0}$ to map them to the real line [dao2022]. We define the transformed parameters as $\boldsymbol{\theta}^{\ast}=\{b^{\ast},A^{\ast},t_{0}^{\ast}\}=\{\log{(b-A)},\log{A},\log{t_{0}}\}$ and the transformation that maps $\boldsymbol{\theta}$ to $\boldsymbol{\theta}^{\ast}$ as $\mathit{T}$.
@@
We will apply stochastic gradient ascent to maximize the ELBO objective function given by Equation (4), using Monte Carlo (MC) estimates of the expectation and AD to calculate gradients with respect to $\boldsymbol{\phi}$ and $\boldsymbol{\zeta}$. However, the gradients of the ELBO with respect to $\boldsymbol{\phi}$ cannot be calculated directly by AD, since the expectation in Equation (4) is taken with respect to $q(\boldsymbol{\theta}^{\ast};\boldsymbol{\phi})$, which depends on $\boldsymbol{\phi}$. We work around this by applying the reparameterization trick [rezende2014, kingma2013], which expresses $\boldsymbol{\theta}^{\ast}$ as a differentiable transformation of an auxiliary noise variable $\boldsymbol{\epsilon}\sim p(\boldsymbol{\epsilon})$, where $p(\boldsymbol{\epsilon})$ does not depend on $\boldsymbol{\phi}$. In particular, we set $p(\boldsymbol{\epsilon})$ to be a standard multivariate Gaussian, $p(\boldsymbol{\epsilon})=\mathcal{N}(\boldsymbol{\epsilon};\mathbf{0},\textbf{I})$, and define the transformation $\widetilde{\boldsymbol{\theta}}^{\ast}=\mathbf{L}\boldsymbol{\epsilon}+\boldsymbol{\mu}$, where $\boldsymbol{\mu}$ and $\mathbf{L}$ are the mean and covariance parameters of the Gaussian approximating density $q(\boldsymbol{\theta}^{\ast};\boldsymbol{\phi})$.
Now we estimate the expectation in Equation (4) with MC samples from $p(\boldsymbol{\epsilon})$. Since the expectation no longer depends on $\boldsymbol{\phi}$, we can use AD directly on these MC estimates to obtain unbiased gradients of the ELBO. We estimate the ELBO for each data point using independent MC samples, since this reduces the variance of the gradients relative to using the same set of MC samples for a given batch of data [kingma2015localreparam]. Thus the ELBO objective for the $i$th data point is estimated by:
@@
VAM inference algorithm and other training details. The VAM training/inference algorithm is summarized in Algorithm 1. The size of the training set was 16,250 samples (65% of the total dataset) for each model. An additional validation set of 3,750 samples (15%) was used to monitor model training. The remaining 5,000 samples (20%) were used to evaluate model performance (the holdout set). We set the batch size $M$ to 256 and the number of MC samples used to estimate the ELBO $L$ to 10. We used the Adam optimizer [kingma2017adam] with the following hyperparameters for model training: learning rate $=1e-3$, $\beta_{1}=0.9$, and $\beta_{2}=0.999$. The same random seed was used to initialize the parameters of all models. Models were trained using a single NVIDIA GeForce RTX 3060 Ti GPU. Each model took approximately one hour to train.
@@
while not converged do
       $\mathbf{t}^{M},\mathbf{c}^{M},\mathbf{s}^{M}\sim\mathbf{x}$ (sample random minibatch of size $M$ from full dataset)
       $\mathbf{v}^{M}\leftarrow\mathrm{CNN}_{\zeta}(\mathbf{s}^{M})$ (calculate drift rate means)
       $\boldsymbol{\epsilon}^{M}\sim\mathcal{N}(\mathbf{0},\textbf{I})$ (draw $M\times L$ samples from the standard multivariate Gaussian)
       $\boldsymbol{\theta}^{\ast M}\leftarrow\mathbf{L}\boldsymbol{\epsilon}^{M}+\boldsymbol{\mu}$ (reparameterize)
       ~$\widetilde{\mathcal{L}}^{M}(\boldsymbol{\phi},\boldsymbol{\zeta})\leftarrow$ calculate ELBO using equation (6)
       $\mathbf{g}^{M}\leftarrow\nabla_{\boldsymbol{\phi},\boldsymbol{\zeta}}\widetilde{\mathcal{L}}^{M}(\boldsymbol{\phi},\boldsymbol{\zeta})$ (calculate gradients using AD)
@@
Convolutional neural network (CNN). We used the same seven-layer CNN architecture for all models (six convolutional layers followed by one fully-connected hidden layer). The number of channels/units used in the seven layers was as follows: 64, 64, 128, 128, 128, 256, 1024. The output layer (after the fully-connected layer) has four channels, one for each drift rate. Each convolutional layer used a 3$\times$3 pixel kernel (stride 1, same padding). Each convolutional layer was followed by a ReLU nonlinearity, instance normalization [ulyanov2017instance], and a max-pooling layer (2$\times$2 pixel window, stride 2), in that order. The fully-connected hidden layer was followed by a ReLU nonlinearity and a dropout layer (dropout rate set to 0.5).
To speed up model training, the first two convolutional layers were initialized with the parameters from a larger CNN trained on an image classification task. Specifically, the pretrained model used a 16-layer VGG architecture and was trained on the ImageNet dataset [deng2009, simonyan2015]. These first two layers were trainable (i.e., not fixed to their initial values). The weights of the other convolutional layers, fully-connected layer and output layer were initialized randomly using the LeCun normal initializer [klambauer2017selfnormalizing]; the biases were initialized with zeros.
@@
We used data augmentation techniques commonly used in other image classification training paradigms to improve the generalization ability of the models. Specifically, for each batch of training data, each image was independently and randomly translated by a small amount. The size of the translation (in pixels) was drawn from a uniform distribution on the interval [0,1] (vertical) and [0,2] (horizontal), rounded to the nearest pixel (horizontal and vertical translations were sampled independently). We also applied a variation of a random elastic image deformation [simard2003elastic] as implemented in the Augmax Python package [augmax]. Specifically, we used the Warp transformation in Augmax and set the strength parameter to 3 and the coarseness parameter to 32. The transformation was applied to each image independently with a probability of 0.75.
@@
The task-optimized CNN models were trained by minimizing the cross-entropy loss function, where the correct label was defined as the true direction of the target bird in each stimulus (i.e., these models were not trained to match the decisions of the human participants, but rather to output the correct target direction). Other than the loss function, all aspects of the training process were the same as those used for the VAM (CNN architecture, optimizer settings, initialization, etc.). For each VAM we trained, we trained one task-optimized model using the same training data ($n$ = 75 task-optimized models).
@@
We analyzed all of the VAMs after 12,800 parameter update steps (corresponding to the 200th training epoch), by which point training had converged. We analyzed all of the task-optimized models after 1,600 parameter update steps (the 25th training epoch) since these models converged much more quickly. All analyses were conducted using a holdout set of $n$ = 5000 LIM stimuli/RTs/choices (i.e., data that was not used to train the models). To generate RTs and choices from the trained models, the LIM stimuli from the holdout set were provided as inputs to the CNN, which output mean drift rates for each stimulus. We denote the drift rate mean for the $k^{th}$ accumulator on trial $i$ as $v_{k}^{(i)}$. We used the LBA model to sample RTs and choices using these mean drift rates, where the drift rate for the $k^{th}$ accumulator on trial $i$, $d_{k}^{(i)}$, is sampled from a normal distribution with mean $v_{k}^{(i)}$ and SD = 1. For a very small fraction of trials (mean $\pm$ s.e.m. percent excluded trials: 0.080 $\pm$ 0.011%, $n$ = 75 models), all of the sampled drift rates were negative, and thus the RT and choice were undefined. These trials were excluded from all analyses.
@@
The models/participants included in the analysis of how RT varied with stimulus layout and horizontal/vertical position were those that exhibited significant modulation of RT by one or more of these stimulus features. Significant modulation was determined by running an ANOVA on the RTs in each stimulus feature bin (e.g., for each layout or each horizontal position bin), using a threshold $p$-value of 0.05. For horizontal stimulus position, we used bins of width = 50 pixels in the original 640$\times$480 pixel window space, except for the leftmost and rightmost bins which had width = 60 pixels. For vertical stimulus position, we used bins of width = 25 pixels.
@@
CNN unit activity. The activation matrices used in the analyses of CNN representations were derived from the responses of units in each layer elicited by the holdout image set. The activations were processed just after the ReLU nonlinearity. For the convolutional layers, we defined the activation of a given unit/channel as the maximum value of that channel calculated across the spatial dimensions [hohman2020summit]. This yields a $N\times K_{l}$ activation matrix, where $N$ is the number of images in the holdout set (5000) and $K_{l}$ is the number of active channels in layer $l$. For the fully-connected layer, the responses of all units were used to define the $N\times K_{l}$ activation matrix directly.
@@
Only incongruent trials were used for all of the analyses involving the CNN activations described below, for the following reasons. For the selectivity and tolerance analyses in which we decoded target or flanker direction from the activity in each layer, described in more detail below, note that the target and flanker directions are by definition the same on congruent trials. As such, a classifier trained to decode target direction from congruent trials could achieve perfect accuracy using information in the unit activity attributable to the flankers, rather than targets. Using only incongruent trials ensured that such cross-contamination could not occur. We used only incongruent trials for all of the other analyses of the activations simply for convenience.
@@
Stimulus feature decoding. To assess how well particular stimulus features could be decoded from the activity in each layer, we trained a linear SVM to classify the values of that stimulus feature using the $N\times K_{l}$ activation matrix in each layer. The activation matrix was standardized before training the classifiers, such that each column had zero mean and unit variance. Horizontal and vertical stimulus positions were discretized to enable classification using the same bins as were used in the behavioral analyses.
@@
The classification task was done in a standard "one-vs-rest" setting: for each value of a given stimulus feature, one sub-classifier was trained on a binary classification task with the chosen value as one class and all other values as the other class, yielding one classifier for each value of the stimulus feature (e.g., four for target direction, seven for stimulus layout). To determine the decision of the combined classifier for a given image, we generated predictions from each sub-classifier and assigned the decision to the sub-classifier with the largest (most confident) prediction. We assessed the overall decoding accuracy of each SVM on a separate test image set. Note that both the training set and test set for the SVMs were subsets of the holdout image set (i.e., the images that were not used to train the VAMs). The SVMs were trained using the LinearSVC model in the scikit-learn Python package [scikit-learn] with the squared hinge loss function and L2 regularization with the penalty parameter C set to 1.0 (the default settings).
@@
Tolerance. To assess the tolerance of model representations to variation in a given stimulus feature (flanker direction, stimulus layout, horizontal position, and vertical position), we trained a linear SVM to classify target direction from the CNN activations using stimuli with the chosen stimulus feature fixed to one value (the training context), and assessed the generalization performance of the SVM on stimuli that contained all other values of that stimulus feature (the generalization context). For example, to assess tolerance to stimulus layout, we trained one SVM to classify target direction using stimuli with the vertical line layout, and assessed the generalization performance of that classifier on stimuli with the six other layouts. We trained one such SVM for each of the seven stimulus layouts, and averaged the generalization performance across these seven SVMs to derive the overall generalization performance measure for a given model and network layer. Other details of the classifiers were the same as those used for the decoding analyses described above.